In 2010, the CPI was 217 and Jim's nominal income was $40,000. In 2011, the CPI was 234 and Jim's nominal income was $42,500. The percentage change in Jim's real income between 2010 and 2011 was approximately:

A.)+ 6.25%

B.)+ 1.5%

C.)- 1.5%

D.)+ 5.9%
Answer #1

Answer

real income =(nominal income /CPI)*100

Real income in 2010=(40000/217)*100=18433.1797

Real income in 2011=(42500/234)*100=18162.3932

the change in real income =((current year real income -last year real income)/last year real income)*100

=((18162.3932-18433.1797)/18433.1797)*100

=-1.46901676

=-1.5%

Option C